OpenStudy (anonymous):

How many liters of pure alcohol must be added to 40 liters of 20% alcohol solution to obtain a mixture which is 25% alcohol?

OpenStudy (amistre64):

a + 40 = t *(-1) a(1) + 40(.2) = t(.25) -------------------- -a +40(-1) = t(-1) a + 40(.2) = t(.25) ------------------- 40(.2-1) = t(.25-1) 40(-.8) ------ = t (-.75) a = t-40 and i prolly messed it up along the way while trying to be coy lol

OpenStudy (amistre64):

prolly could have gone the easier route and *-.25 on top to just solve for a :)

OpenStudy (anonymous):

8/3 liters?

OpenStudy (amistre64):

a + 40 = t *(-.25) a(1) + 40(.2) = t(.25) -------------------- a(-.25) +40(-.25) = t(-.25) a (1) + 40(.2) = t(.25) ----------------------- a(1-.25) +40(.20-.25)=0 a(.75) =40(.25-.20) 40(.25-.20) a = ----------- .75

OpenStudy (amistre64):

8/3 is what i get too :)

OpenStudy (anonymous):

One can simply solve for \( x \) in \[ (8+x) =(40+x) \times 0.25 \]

OpenStudy (amistre64):

\[\frac{8}{3}=2+\cfrac{2}{3}\to 2+\cfrac{1}{\cfrac{3}{2}}\to 2+\cfrac{1}{1+\cfrac{1}{2}} \] \[[2;1,2]\]as a continued fraction :)

OpenStudy (anonymous):

Yeah sure, @amistre64 is known for his cryptic approaches lol

OpenStudy (anonymous):

100% L + 20% 40 = 25% ( L + 40) -> L + 8 = L/4 + 10 -> 3L/4 = 2 => L = 8/3 litter pure alcohol

OpenStudy (amistre64):

kitty litter ... ewww!!!

OpenStudy (anonymous):

Hey, I guess I'm obsessed with my kitty :")